# The norms for symmetric and antisymmetric tensor products of the weighted shift operators

### v1

@@ 3.1 Proof of “$(1)\Leftrightarrow(2)$
$\sum\limits_{t=1}^{\mathscr{M}-1}\sum\limits_{\left\langle\mathfrak{j}\right\rangle\in E_{k+S_{\mathfrak{l}},t}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right)e_{\langle\mathfrak{j}\rangle}$
@@
$\left\|\left(S_{\alpha,{l_{1}}}\odot S_{\alpha,{l_{2}}}\odot\cdots\odot S_{\alpha,{l_{n}}}\right)\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}a_{\langle\mathfrak{i}\rangle}\frac{e_{\langle\mathfrak{i}\rangle}}{\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right\|^{2}$ (3.7)
$\geq$ $\left\|\sum\limits_{\left\langle\mathfrak{j}\right\rangle\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right)e_{\langle\mathfrak{j}\rangle}\right\|^{2}$
$=$ $\sum\limits_{\left\langle\mathfrak{j}\right\rangle\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left|\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right|^{2}\left\|e_{\langle\mathfrak{j}\rangle}\right\|^{2}$
$=$ $\frac{1}{n!}\sum\limits_{\left\langle\mathfrak{j}\right\rangle\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left|\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right|^{2}$
$\geq$ $\frac{{}^{\#}\tilde{A}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}}{P(k,n)}\left(1-\varepsilon\right)^{2\left(\left|l_{1}\right|+\cdots+\left|l_{n}\right|\right)},$
where the last inequality comes from Lemma 2.4 and Lemma 2.8, together with the fact that $\left\langle\mathfrak{i}\right\rangle\in R_{\mathfrak{j}}$ if and only if $N_{\mathfrak{j},\mathfrak{i}}\neq\emptyset.$ By Lemma 2.7
$\lim\limits_{k\to\infty}\frac{{}^{\#}\tilde{A}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}}{P(k,n)}\left(1-\varepsilon\right)^{2\left(\left|l_{1}\right|+\cdots+\left|l_{n}\right|\right)}=\left(1-\varepsilon\right)^{2\left(\left|l_{1}\right|+\cdots+\left|l_{n}\right|\right)},$
@@
by (3.8) and (3.9), it suffices to prove that
@@
$\begin{cases}|\alpha_{i_{j}}|=1,&j=1,\cdots,N,\\ |\alpha_{i}|<\delta,&i\not\in\{i_{j};j=1,\cdots,N\}.\end{cases}$
@@
In fact, for $(i_{1},i_{2})\in\mathbb{Z}^{2}$, let $a_{\left\langle i_{1},i_{2}\right\rangle}\in\mathbb{C}$ such that $a_{\left\langle i_{1},i_{2}\right\rangle}=0$ for $(i_{1},i_{2})\notin\mathbb{N}^{2}$. By Lemma 2.1, for every $k\geq 0,$
$\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{\langle i_{1},i_{2}\rangle}{\left(e_{i_{1}}\odot e_{i_{2}}\right)}=\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{\langle i_{1},i_{2}\rangle}\beta_{i_{1},K}\beta_{i_{2},K}\left(e_{i_{1}+K}\odot e_{i_{2}+K}\right).$
@@
$\left\|\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{\langle i_{1},i_{2}\rangle}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|}\right\|^{2}$ (3.14)
$=\sum\limits_{0\leq j_{1}\leq j_{2}\atop j_{1}+j_{2}=k+2K}\left|\frac{a_{\left\langle j_{1}-K,j_{2}-K\right\rangle}\beta_{j_{1}-K,K}\beta_{j_{2}-K,K}}{\left\|e_{j_{1}-K}\odot e_{j_{2}-K}\right\|}\right|^{2}\left\|e_{j_{1}}\odot e_{j_{2}}\right\|^{2}.$
Thus when $\sum\limits_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}|a_{\langle i_{1},i_{2}\rangle}|^{2}=1,$ by (3.14),
$\left\|\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{\langle i_{1},i_{2}\rangle}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|}\right\|^{2}$ (3.15)
$\leq\delta^{2}\left(\sum\limits_{0\leq j_{1}\leq j_{2}\atop j_{1}+j_{2}=k+2K}\left|\frac{a_{\left\langle j_{1}-K,j_{2}-K\right\rangle}}{\left\|e_{j_{1}-K}\odot e_{j_{2}-K}\right\|}\right|^{2}\left\|e_{j_{1}}\odot e_{j_{2}}\right\|^{2}\right)$
$\leq\delta^{2}\left(\sum\limits_{0\leq j_{1}\leq j_{2}\atop j_{1}+j_{2}=k+2K}\left|a_{\left\langle j_{1}-K,j_{2}-K\right\rangle}\right|^{2}\right)\leq\delta^{2}\left(\sum\limits_{0\leq j_{1}\leq j_{2}\atop j_{1}+j_{2}=k}\left|a_{\left\langle j_{1},j_{2}\right\rangle}\right|^{2}\right)=\delta^{2}<1.$
@@
$\left\langle\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k_{1}}a_{\langle i_{1},i_{2}\rangle}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|},\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k_{2}}a_{\langle i_{1},i_{2}\rangle}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|}\right\rangle=0,$
@@
$\left\|\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\right\|=\sup\limits_{k\in\mathbb{N}}\left\{\left\|\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{\langle i_{1},i_{2}\rangle}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|}\right\|\right\}\leq\delta.$
$\square$
3.2 Proof of “$(1)\Leftrightarrow(3)$
In this subsection, we prove $(1)\Leftrightarrow(3)$. First, we need some notations and some lemmas. It will be seen that the lemmas listed below are similar to those in Section 2, however we need more techniques from antisymmetric products to deal with the proofs.
@@ 3.2 Proof of “$(1)\Leftrightarrow(3)$
$W^{\prime}=\{\mathfrak{i}\in W:i_{r}-i_{r-1}>4|\mathfrak{l}|~{}for~{}all~{}2\leq r\leq n\}.$
@@
For $\mathfrak{j}^{\prime}\in W^{\prime},\mathfrak{j}\in\mathfrak{W}\backslash W^{\prime},$ we have
@@
Notice that $j_{t}^{\prime}-j_{t-1}^{\prime}>4|\mathfrak{l}|$ for all $2\leq t\leq n.$ Hence (3.17) holds.
@@
For $\mathfrak{j}\in\mathfrak{W}\backslash W,$ by the proof of Lemma 3.2, it can be shown easily that there exists a constant $C_{1}^{\prime}>0$, such that
$\mathfrak{W}_{(k,n)}-W_{(k,n)}\leq C_{1}^{\prime}k^{n-2}.$ (3.20)
For $\mathfrak{j}\in W\backslash W^{\prime},$ by the definitions of $W$ and $W^{\prime},$ there exists an integer $2\leq r_{1}\leq n$ such that $0<j_{r_{1}-1}-j_{r_{1}}\leq 4|\mathfrak{l}|,$ which implies that there exists a constant $C_{2}^{\prime}>0$, such that
$W_{(k,n)}-W^{\prime}_{(k,n)}\leq C_{2}^{\prime}k^{n-2}.$ (3.21)
@@
#$^{\#}\mathcal{R}_{\mathfrak{i}}\leq\ ^{\#}\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\ \right\}=\mathscr{M}.$ (3.23)
@@
$\mathcal{A}_{k}\backslash\mathcal{A}_{k}^{\prime}=\left\{\mathfrak{i}\in\mathfrak{W}\backslash W^{\prime}:|\mathfrak{i}|=k\ and\ ^{\#}\mathcal{R}_{\mathfrak{i}}<\mathscr{M}\right\}.$
Hence by (3.19), there exists a constant $C_{1}>0$ such that
#$^{\#}\mathcal{A}_{k}\leq^{\#}\mathcal{A}_{k}^{\prime}+C_{1}k^{n-2}.$ (3.26)
@@
$R_{\mathfrak{i}}\subseteq\{\left\langle\mathfrak{i}-\mathfrak{l}_{\pi}\right\rangle\in\mathbb{N}^{n}/\Sigma_{n}:\pi\in\Sigma_{n}\}.$ (3.28)
Since $\mathfrak{i}\in\mathcal{A}_{k}^{\prime}\subseteq W^{\prime}\subseteq W,$ we have $i_{r}-i_{r-1}>4|\mathfrak{l}|$ for all $2\leq r\leq n,$ hence $d\leq i_{1}-l_{\pi(1)}<\cdots<i_{n}-l_{\pi(n)}.$ Then by (3.22), it is easy to see that
$\{\mathfrak{i}-\mathfrak{l}_{\pi}\in\mathbb{N}^{n}:\pi\in\Sigma_{n}\}\subseteq\mathcal{R}_{\mathfrak{i}}$ (3.29)
Therefore by (3.28) and (3.29), we have ${}^{\#}R_{\mathfrak{i}}\leq^{\#}\mathcal{R}_{\mathfrak{i}},$ which implies that (3.27) holds. The claim is proved. By (2.14), (2.15) and (2.16), there exists a constant $C_{2}>0$, such that ${}^{\#}A_{k}^{\prime}\leq C_{2}k^{n-2}.$ Hence (3.26) and (3.27) ensure that there exists a constant $M>0$, such that ${}^{\#}\mathcal{A}_{k}\leq Mk^{n-2}$.
@@
${}^{\#}\mathcal{N}_{\mathfrak{j},\mathfrak{i}}$ =#$=\ ^{\#}\left\{\pi\in\Sigma_{n}:e_{i_{1}+l_{\pi(1)}}\wedge\cdots\wedge e_{i_{n}+l_{\pi(n)}}=e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}\right\}\geq\ n_{1}!\cdots n_{k}!.$
@@
#$^{\#}\check{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\leq Ck^{n-2}.$ (3.36)
@@
and for $10<\varepsilon<$ 1, set
$\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}=\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\cap\mathcal{E}_{k+S_{\mathfrak{l}},\mathscr{M}}^{\prime}.$ (3.38)
@@
$\bigcup_{j=0}^{\mathscr{M}}\mathcal{E}_{k+S_{\mathfrak{l}},j}=\left\{\mathfrak{j}\in\mathfrak{W}_{T}:\left|\mathfrak{j}\right|=k+S_{\mathfrak{l}}\right\}.$ (3.39)
@@
${}^{\#}\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}$ =#$=\ ^{\#}\mathcal{E}_{k+S_{\mathfrak{l}},\mathscr{M}}-\ ^{\#}\check{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}$ (3.40)
@@
In fact, for every $10\leq j\leq\mathscr{M}-$ 1, by (3.30), (3.24) and Lemma 3.4,
#$^{\#}\mathcal{E}_{k+S_{\mathfrak{l}},j}\leq\ ^{\#}\mathcal{A}_{k+S_{\mathfrak{l}}}\leq M\left(k+S_{\mathfrak{l}}\right)^{n-2}.$ (3.42)
@@
Hence from (2.44), (2.45), (3.43), (3.42) and Lemma 3.5, for every $10\leq j\leq\mathscr{M}-$ 1,
@@
$\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\backslash\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}=\left\{\mathfrak{j}\in\mathcal{E}_{k+S_{\mathfrak{l}},\mathscr{M}}\backslash\mathcal{E}_{k+S_{\mathfrak{l}},\mathscr{M}}^{\prime}:\Gamma_{j_{m},-l_{\pi(m)}}>1-\varepsilon,\forall 1\leq m\leq n,\forall\,\pi\in\Sigma_{n}\right\}$
@@
$\mathcal{P}_{\mathfrak{i},k,t}=\mathop{\bigcup}\limits_{\mathfrak{j}\in\mathcal{E}_{k+S_{\mathfrak{l}},t}}\mathcal{N}_{\mathfrak{j},\mathfrak{i}}=\left(\mathop{\bigcup}\limits_{\mathfrak{j}\in\mathcal{E}_{k+S_{\mathfrak{l}},t}\backslash\mathcal{E}_{k+S_{\mathfrak{l}},t}^{\prime}}\mathcal{N}_{\mathfrak{j},\mathfrak{i}}\right)\bigsqcup\left(\bigsqcup\limits_{\mathfrak{j}\in\mathcal{E}_{k+S_{\mathfrak{l}},t}^{\prime}}\mathcal{N}_{\mathfrak{j},\mathfrak{i}}\right).$ (3.50)

### v2

@@ 3.1 Proof of $(1)\Leftrightarrow(2)$
$\sum\limits_{t=1}^{\mathscr{M}-1}\sum\limits_{[\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},t}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right)e_{[\mathfrak{j}]}$
@@
$\left\|\left(S_{\alpha,{l_{1}}}\odot S_{\alpha,{l_{2}}}\odot\cdots\odot S_{\alpha,{l_{n}}}\right)\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}a_{[\mathfrak{i}]}\frac{e_{[\mathfrak{i}]}}{\left\|e_{[\mathfrak{i}]}\right\|}\right\|^{2}$ (3.7)
$\geq$ $\left\|\sum\limits_{[\mathfrak{j}]\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right)e_{[\mathfrak{j}]}\right\|^{2}$
$=$ $\sum\limits_{[\mathfrak{j}]\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left|\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right|^{2}\left\|e_{[\mathfrak{j}]}\right\|^{2}$
$=$ $\frac{1}{n!}\sum\limits_{[\mathfrak{j}]\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left|\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right|^{2}$
$\geq$ $\frac{{}^{\#}\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}{P(k,n)}\left(1-\varepsilon\right)^{2\left(\left|l_{1}\right|+\cdots+\left|l_{n}\right|\right)},$
where the last inequality comes from Lemma 2.4 and Lemma 2.8, together with the fact that $\mathfrak{i}]\in R_{\mathfrak{j}}$ if and only if $N_{\mathfrak{j},\mathfrak{i}}\neq\emptyset.$ By Lemma 2.7
$\lim\limits_{k\to\infty}\frac{{}^{\#}\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}{P(k,n)}\left(1-\varepsilon\right)^{2\left(\left|l_{1}\right|+\cdots+\left|l_{n}\right|\right)}=\left(1-\varepsilon\right)^{2\left(\left|l_{1}\right|+\cdots+\left|l_{n}\right|\right)},$
@@
by (3.8) and (3.9), it suffices to prove that for $i\geq 0,$
@@
$\begin{cases}|\alpha_{i_{j}}|=1,&j=1,\cdots,N,\\ |\alpha_{i}|<\delta,&i\not\in\{i_{j}:j=1,\cdots,N\}.\end{cases}$
@@
In fact, for $(i_{1},i_{2})\in\mathbb{Z}^{2}$, let $a_{[i_{1},i_{2}]}\in\mathbb{C}$ such that $a_{[i_{1},i_{2}]}=0$ for $(i_{1},i_{2})\notin\mathbb{N}^{2}$. By Lemma 2.1, for every $k\geq 0,$
$\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{[i_{1},i_{2}]}{\left(e_{i_{1}}\odot e_{i_{2}}\right)}=\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{[i_{1},i_{2}]}\beta_{i_{1},K}\beta_{i_{2},K}\left(e_{i_{1}+K}\odot e_{i_{2}+K}\right).$
@@
$\left\|\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{[i_{1},i_{2}]}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|}\right\|^{2}$ (3.14)
$\leq\sum\limits_{0\leq j_{1}\leq j_{2}\atop j_{1}+j_{2}=k+2K}\left|\frac{a_{[j_{1}-K,j_{2}-K]}\beta_{j_{1}-K,K}\beta_{j_{2}-K,K}}{\left\|e_{j_{1}-K}\odot e_{j_{2}-K}\right\|}\right|^{2}\left\|e_{j_{1}}\odot e_{j_{2}}\right\|^{2}.$
Thus when $\sum\limits_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}|a_{[i_{1},i_{2}]}|^{2}=1,$ by (3.14),
$\left\|\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{[i_{1},i_{2}]}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|}\right\|^{2}$ (3.15)
$\leq\delta^{2}\left(\sum\limits_{0\leq j_{1}\leq j_{2}\atop j_{1}+j_{2}=k+2K}\left|\frac{a_{[j_{1}-K,j_{2}-K]}}{\left\|e_{j_{1}-K}\odot e_{j_{2}-K}\right\|}\right|^{2}\left\|e_{j_{1}}\odot e_{j_{2}}\right\|^{2}\right)$
$\leq\delta^{2}\left(\sum\limits_{0\leq j_{1}\leq j_{2}\atop j_{1}+j_{2}=k+2K}\left|a_{[j_{1}-K,j_{2}-K]}\right|^{2}\right)\leq\delta^{2}\left(\sum\limits_{0\leq j_{1}\leq j_{2}\atop j_{1}+j_{2}=k}\left|a_{\left\langle j_{1},j_{2}\right\rangle}\right|^{2}\right)=\delta^{2}<1.$
@@
$\left\langle\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k_{1}}a_{[i_{1},i_{2}]}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|},\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k_{2}}a_{[i_{1},i_{2}]}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|}\right\rangle=0,$
@@
$\left\|\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\right\|=\sup\limits_{k\in\mathbb{N}}\left\{\left\|\left(S_{\alpha,{K}}\odot S_{\alpha,{K}}\right)\sum_{0\leq i_{1}\leq i_{2}\atop i_{1}+i_{2}=k}a_{[i_{1},i_{2}]}\frac{e_{i_{1}}\odot e_{i_{2}}}{\left\|e_{i_{1}}\odot e_{i_{2}}\right\|}\right\|\right\}\leq\delta.$
3.2 Proof of $(1)\Leftrightarrow(3)$
In this subsection, we prove $(1)\Leftrightarrow(3)$. First, we need some notations and some lemmas. It will be seen that the lemmas listed below are similar to those in Section 2, however we need more techniques from antisymmetric tensor products to deal with the proofs.
@@ 3.2 Proof of $(1)\Leftrightarrow(3)$
$W^{\prime}=\{\mathfrak{i}\in W:i_{1}\geq d+|\mathfrak{l}|,i_{r}-i_{r-1}>4|\mathfrak{l}|~for~all~2\leq r\leq n\}.$
@@
[…]
@@
Notice that $j_{t}^{\prime}-j_{t-1}^{\prime}>4|\mathfrak{l}|$ for all $2\leq t\leq n.$ Hence (3.17) holds. ∎
@@
For $\W,\mathfrak{j}\in\mathfrak{W}\backslash$ W, by the proof of Lemma 3.2, it can be shown easily that there exists a constant $C_{1}^{\prime}>0$, such that
$\mathfrak{W}_{(k,n)}-W_{(k,n)}\leq C_{1}^{\prime}\left(k+1\right)^{n-2}.$ (3.20)
[…]
$W_{(k,n)}-W^{\prime}_{(k,n)}\leq C_{2}^{\prime}\left(k+1\right)^{n-2}.$ (3.21)
@@
$^{\#}\mathcal{R}_{\mathfrak{i}}\leq\ ^{\#}\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\ \right\}=\mathscr{M}.$ (3.23)
@@
[…]
Hence by (3.19), there exists a constant $C_{1}>0$ such that for $k>0$ sufficiently large,
$^{\#}\mathcal{A}_{k}\leq^{\#}\mathcal{A}_{k}^{\prime}+C_{1}k^{n-2}.$ (3.26)
@@
$R_{\mathfrak{i}}\subseteq\{[\mathfrak{i}-\mathfrak{l}_{\pi}]\in\mathbb{N}^{n}/\Sigma_{n}:\pi\in\Sigma_{n}\}.$ (3.28)
Since $\mathfrak{i}\in\mathcal{A}_{k}^{\prime}\subseteq W^{\prime}\subseteq W,$ we have $i_{r}-i_{r-1}>4|\mathfrak{l}|$ for all $2\leq r\leq n$ and $i_{1}\geq d+|\mathfrak{l}|$, hence $d\leq i_{1}-l_{\pi(1)}<\cdots<i_{n}-l_{\pi(n)}$ for every $\pi\in\Sigma_{n}.$ Then by (3.22), it is easy to see that
$\{\mathfrak{i}-\mathfrak{l}_{\pi}\in\mathbb{N}^{n}:\pi\in\Sigma_{n}\}\subseteq\mathcal{R}_{\mathfrak{i}}.$ (3.29)
Therefore by (3.28) and (3.29), we have $^{\#}R_{\mathfrak{i}}\leq^{\#}\mathcal{R}_{\mathfrak{i}},$ which implies that (3.27) holds. The claim is proved. By (2.14), (2.15) and (2.16), there exists a constant $C_{2}>0$, such that $^{\#}A_{k}^{\prime}\leq C_{2}\left(k+1\right)^{n-2}.$ Hence (3.26) and (3.27) ensure that for $k>0$ sufficiently large, there exists a constant $M>0$, such that $^{\#}\mathcal{A}_{k}\leq Mk^{n-2}$. ∎
@@
${}^{\#}\mathcal{N}_{\mathfrak{j},\mathfrak{i}}$ $=\ ^{\#}\left\{\pi\in\Sigma_{n}:e_{i_{1}+l_{\pi(1)}}\wedge\cdots\wedge e_{i_{n}+l_{\pi(n)}}=e_{j_{1}}\wedge\cdots\wedge e_{j_{n}}\right\}\geq\ n_{1}!\cdots n_{k}!.$
@@
$^{\#}\check{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\leq Ck^{n-2}.$ (3.36)
@@
and for $0<\varepsilon<1$, set
$\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}=\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\cap\mathcal{E}_{k+S_{\mathfrak{l}},\mathscr{M}}^{\prime}$ (3.38)
and
$\check{\mathcal{A}}^{\prime}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}=\check{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\cap\mathcal{E}_{k+S_{\mathfrak{l}},\mathscr{M}}^{\prime}.$
@@
$\bigcup_{j=0}^{\mathscr{M}}\mathcal{E}_{k+S_{\mathfrak{l}},j}=\left\{\mathfrak{j}\in\mathfrak{W}:\left|\mathfrak{j}\right|=k+S_{\mathfrak{l}}\right\}.$ (3.39)
@@
${}^{\#}\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}$ $=\ ^{\#}\mathcal{E}_{k+S_{\mathfrak{l}},\mathscr{M}}-\ ^{\#}\check{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}$ (3.40)
@@
In fact, for every $0\leq j\leq\mathscr{M}-1$, by (3.30), (3.24) and Lemma 3.4,
$^{\#}\mathcal{E}_{k+S_{\mathfrak{l}},j}\leq\ ^{\#}\mathcal{A}_{k+S_{\mathfrak{l}}}\leq M\left(k+S_{\mathfrak{l}}\right)^{n-2}.$ (3.42)
@@
Hence from (2.44), (2.45), (3.43), (3.42) and Lemma 3.5, for every $0\leq j\leq\mathscr{M}-1$,
@@
[…]
@@
$\lim\limits_{k\to\infty}\frac{{}^{\#}\tilde{\mathcal{A}}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}}{W_{(k,n)}}=1.\qed$
@@
[…]